OpenStudy (anonymous):

x/5 + 9 = 4 who know how to so this ??

OpenStudy (anonymous):

We need to find x so we have to put x on one side and the other on the other side x/5 + 9 = 4 x/5 +9 -9 = 4 -9 x/5=-5 x/5 (5) = -5(5) x=-25

OpenStudy (anonymous):